OpenStudy (anonymous):

There are 25 desks in a classroom. The number of rows is five less than twice the number of desks in each row. Find the number of rows and the number of desks in each row.

OpenStudy (anonymous):

if you think of every desk as a unit of the classroom than you get a box/rectangle of desks/units, so you end up with the Area of a rectangle equation \[A=LW\] \[25=LW\]

OpenStudy (mathstudent55):

Let r = number of rows Let d = number of desks per row "The number of rows is five less than twice the number of desks in each row." r = 2d - 5 "There are 25 desks in a classroom." The number of desks, 25, is the number of rows multiplied by the number of desks per row, so rd = 25 Since we have that r = 2d - 5, substitute: (2d - 5)d = 25 Solve the equation for d

OpenStudy (anonymous):

\[2d ^{2}-5d-25=0\] ??

OpenStudy (mathstudent55):

Correct. Now solve the equation.

OpenStudy (anonymous):

so d=-5/2 and d=5 throw away the negative value and you have d=5

OpenStudy (anonymous):

that would equal the number of desks per row, what is the equation for number of rows?

OpenStudy (anonymous):

oh wait, D=5 and R=5 ??

OpenStudy (mathstudent55):

correct

OpenStudy (anonymous):

thanks!